OpenStudy (anonymous):

Can someone help me with an Algebra 1 question?

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)?

OpenStudy (anonymous):

y = 3x + 6

OpenStudy (anonymous):

Not one of my options :/

OpenStudy (anonymous):

do you need me to show the work?

OpenStudy (anonymous):

Sure, but its not one of the answers.

OpenStudy (anonymous):

10 - 1 / 1 +2 = 9 /3 = 3 , and that's the slope of the equation. then u substitute any of the x y coordinate above into the equation to find y-intercept, in this case we'll use (1,10) 10 = 1x3 + b 10 = 3 + b b= 7 sorry, the equation should be y = 3x + 7 or in standard form 3x -y = -7 or 3x - 7 + 7 = 0

OpenStudy (anonymous):

There you go, "3x-y=-7" is one of my options. Thank you!

OpenStudy (anonymous):

No Problem